OpenStudy (geometry_hater):

Which is an equation of the line passing through points at (1, 6) and (-2, 1)?

OpenStudy (cathyangs):

Find the slope of the line, m. then using the slope and one of the points in point-slope form, which you can simplify to slope-intercept.

OpenStudy (cathyangs):

*then use the slope and one of the points in point-slope form,

OpenStudy (slaaibak):

First get the slope. \[{y_2 - y_1} \over {x_2 - x_1}\] Then use this formula with any of the coordinate pairs \[y - y_1 = m(x-x_1)\]

OpenStudy (geometry_hater):

so the answer is what? y = -1/2x +13/2?

OpenStudy (anonymous):

y=mx+b Gradient= m = y1-y2/x1-x2 = 6-1/1+2 4/3 y=4/3x+b Then replace (1,6) in (x,y) = 6=4/3*1+b 14/3=b y=4/3x+14/3 if all calculations are correct

OpenStudy (anonymous):

sorry, 5/3 my calculations are wrong

OpenStudy (anonymous):

the principal is the same, however... replace 4 with 5, and redo it, and you should get the answer...

OpenStudy (anonymous):

y=5/3x+13/3 should be the answer now

OpenStudy (anonymous):

Start with the basic eqn of a line: y=mx+b You can figure out the slope of the line by calculating\[(y_2-y_1)/(x_2-x_1)\] Which is 5/3. Then you can plug any of the two points into the line equation to find b: 6=(5/3)(1) + b Solve for b, which is 13/3 and you get the equation y=5/3 x + 13/3